OpenStudy (anonymous):

Electron exhibiting a particle wavelength of 9.88 x 10-9 are emitted from hydrogen atoms after interaction with high energy photons. Assuming that the electrons in the hydrogen atom were in their ground state prior to interaction, determine the wavelength of the photon which generated the free electrons

OpenStudy (aaronq):

find the energy necessary to ionize the electron using the rydberg formula. find the kinetic energy of the electron: use the deBroglie equation and \(KE=\dfrac{1}{2}*mv^2\) add these 2 energies and convert to wavelength using planck's relation.

OpenStudy (anonymous):

1. Rydberg's equation has only integers for n (initial, final) and a constant and a wavelength value. There is no energy in that equation. 2. How will we find the velocity when even the energy is unknown?

OpenStudy (aaronq):

yes it does, \(E=\dfrac{1}{\lambda}=R_0(\dfrac{1}{n_f^2}-dfrac{1}{n_i^2}\)

OpenStudy (aaronq):

\(E=\dfrac{1}{\lambda}=R_0(....)\)

OpenStudy (anonymous):

okay got it, but how does one proceed to the next step of finding the kinetic energy? cant find the velocity with two unknowns right?

OpenStudy (aaronq):

de broglie equation: \(\lamda =\dfrac{h}{mv}\) find v. enter into KE=1/2mv^2

OpenStudy (aaronq):

\(\lambda =\dfrac{h}{mv}\)

OpenStudy (anonymous):

I understood this de broglie one but I must say you are mistaken regarding the rydberg formula. 1/lambda is equal to the wave number and not the energy sir.

OpenStudy (aaronq):

sorry, you're right, you gotta use plancks relation again.

OpenStudy (anonymous):

wait, so which is the set of equations that must be used then?